OpenStudy (anonymous):

Perform the indicated computation. Write the answer in scientific notation. 12 x 10^4/ 3 x 10 ^-2 A. 8 x 102 B. 4 x 106 C. 4 x 102 D. 8 x 106

OpenStudy (anonymous):

2 . Find the product. (7x - 6)2 A. 49x2 + 36 B. 49x2 - 84x + 36 C. 7x2 - 84x + 36 D. 7x2 + 36

OpenStudy (anonymous):

For the second one, u distribute the ^2 so you would have 49x^2 - 36

OpenStudy (agent0smith):

12 x 10^4/ 3 x 10 ^-2 \[\Large \frac{ 12 \times 10 ^{4} }{ 3 \times 10 ^ {-2} } = \frac{ 12 }{ 3 } \times \frac{ 10 ^{4} }{ 10 ^ {-2} }\]

OpenStudy (agent0smith):

use \[\Large \frac{ x^a }{ x^b } = x^{a-b}\] so the \[\Large \frac{ 10 ^{4} }{ 10 ^ {-2} } = 10^{4- (-2)}\] (that is not the final answer)

OpenStudy (agent0smith):

@Rexy705 that's not quite correct. \[\Large (7x - 6)^2 = (7x- 6) (7x-6)\]now simplify that

OpenStudy (anonymous):

oh that's right that totally slipped my mind =P

OpenStudy (anonymous):

And then u would use foil

OpenStudy (anonymous):

could you help me with another please @agent0smith ???

OpenStudy (anonymous):

add or subtract  a. 2x - 16 / (x+2) ( x-2) b. 2/(x+2) (x-2) c . 2x-4/(x+2) ( x-2) d. 2x+16//(x+2)(x-2)